# ClaimDB: A Fact Verification Benchmark over Large Structured Data

### v1

@@ 7 Conclusion
In this work, we introduce ClaimDB, a fact-verification benchmark over large-scale structured data. It contains 80 unique databases, each with an average of 11 tables and 4.5M records. Each claim is grounded in evidence deliberately composed of millions of records across multiple tables. ClaimDB is a clear step toward real-world claim verification—mirroring the various challenges of high-stakes verification in our modern world.
@@ 8 Limitations
Reliance on BIRD. Our benchmark is derived from NL-to-SQL pairs in the BIRD benchmark (Section 3.2). As a result, any annotation error in BIRD directly propagates to ClaimDB. For example, if the NL question is “How old was the oldest mayor of Seattle when elected?” but the associated SQL query instead computes the answer for Portland, any claims generated from this pair would be incorrectly labeled. To mitigate this risk for public evaluation and model development, we construct our test splits exclusively from claims derived from the dev split of BIRD, which has undergone substantial subsequent cleanup by prior work (Section 5). Importantly, our dependency on BIRD is not irreversible: as additional errors in BIRD are identified, the corresponding derived claims can be removed from ClaimDB on-the-fly.
Single Evidence Modality. ClaimDB deliberately focuses on structured data as its primary evidence source (which may include numerical, categorical, temporal, and short textual fields). While real-world fact verification often involves multi-modal evidence—such as free-form text, reports, news articles, charts, and other unstructured sources—structured data remains a critical yet comparatively underexplored modality in existing benchmarks. By isolating this setting, ClaimDB enables targeted evaluation of reasoning over complex structured evidence. We acknowledge that this focus alone does not capture the full breadth of real-world verification, and view extension to multi-modal evidence as a natural direction for future work.
Snapshot Validity and Evaluation Scope. Entailment, contradiction, and NEI are assessed with respect to the fixed database snapshot used to construct the benchmark. Because the underlying databases come from established public dataset releases—and are therefore static and potentially dated—answers may differ if a system incorporates newer external information. For example, a query about the “bottom-5 enrollment schools” could change, even if the claim was correct in the snapshot used for claim generation. As with many ML benchmarks, evaluation should therefore be interpreted relative to the provided database state Li et al. (2023) rather than a live, continuously updated world.
@@ 9 Ethical Considerations
Data Licensing. ClaimDB bases all claims on NL-to-SQL pairs coming from the BIRD benchmark which is released under CC BY-SA 4.0 license. ClaimDB is released under the same license, in accordance with the original terms.
@@ 10 Acknowledgments
This work was supported by NSF III 2507117, NSF SHF 2312195, and NSF IIS 2314527. We also thank David Alexander for an insightful discussion on benchmark performance.
Appendix A Evaluation
A.1 Evaluation Setup
@@
LLM Backbone Macro-F1 Acc. Tokens
gpt-5-nano (2025b) 0.791 0.793 20M
gpt-5-mini (2025b) 0.828 0.828 19M
gpt-oss:20b (2025) 0.759 0.763 14M
gemini-2.5-flash (2025) 0.761 0.758 8M
gemini-3-flash (2025) 0.807 0.805 12M
claude-3-5-haiku (2024) 0.671 0.673 11M
claude-haiku-4-5 (2025) 0.799 0.792 35M
qwen3:1.7b (2025) 0.231 0.363 2M
qwen3:4b (2025) 0.461 0.493 11M
qwen3:8b (2025) 0.493 0.529 10M
qwen3:14b (2025) 0.465 0.506 9M
qwen3:32b (2025) 0.539 0.558 9M
qwen3-coder:30b (2025) 0.685 0.686 22M
qwq:32b (2025) 0.498 0.513 8M
mistral-nemo:12b (2024) 0.345 0.354 8M
ministral-3:3b (2025b) 0.332 0.369 15M
ministral-3:8b (2025b) 0.543 0.556 16M
ministral-3:14b (2025b) 0.618 0.618 23M
mistral-small:22b (2025c) 0.296 0.379 15M
magistral:24b (2025a) 0.430 0.468 18M
devstral:24b (2025a) 0.388 0.420 -
devstral-small-2 (2025b) 0.572 0.576 16M
nemotron-3-nano (2025) 0.660 0.656 30M
llama3.1:8b (2024) 0.301 0.354 7M
llama3.2:3b (2024) 0.317 0.358 8M
cogito:14b (2025) 0.384 0.450 8M
cogito:32b (2025) 0.548 0.567 11M
Table 3: Private test set results. We report macro-F1 and accuracy. We also report the total number of tokens (input + output) across the 1,000 experiments for each model. Per-label statistics are omitted to preserve the integrity of the private test set and evaluation.
@@
ministral-3:3b 256K 0.15
ministral-3:8b 256K 0.15
ministral-3:14b 256K 0.15
@@
A.2 Private Test Set
The experimental results on the private test set are shown in Table 3. To reduce the risk of test-set exploitation and benchmark gaming, we report only aggregate metrics (Macro-F1 and accuracy), omitting per-label and fine-grained analyses.
A.3 Structured Outputs
@@
For example, some leading agentic frameworks simulate structured outputs by injecting a prompt that forces the model to call a predefined tool for its final answer [^3] https://github.com/pydantic/pydantic-ai/issues/242 , [^4] https://github.com/openai/openai-agents-python/issues/1778#issuecomment-3316092585, rather than enforcing the schema at decoding time when the GCD option is available (e.g., on open-source models). Aside from bloating the model context, this approach is inherently brittle: the LLM can simply ignore the instruction.
@@ Appendix C Prompts
In Figures 12, 13 and 14 we showcase the prompts used in creating the claims (Section 3.5). In Figures 15 and 16 we illustrate the two prompts used for the judging process—one for contradicted and entailed claims and one for NEI since the latter has extra rubrics and extra golden context (Section 4.1). Lastly, in Figure 17, we show the single, optimized verifier prompt for all models in our evaluations.
@@
Figure 17: Prompt for all verifier models in the experiments (Section 6.1). It has been carefully constructed with many iterations and refinements across the different model families. We have tried our best to make the models perform as good as possible. For example, after noticing some models deciding not to use tools and hallucinate an answer, we including instructions like “Use the available tools to query the database and gather evidence before making a decision” and “You should always start by querying the database for the schema”. Furthermore, for a discussion on output format see Appendix A.3. Notably, for the observant readers, there is no explicit explanation of the tool the agent has in its disposal because these descriptions are provided by the frameworks internally (most are built this way) by injecting the docstrings and input-output schemas of the tools in the prompt using templates.

### v2

@@ 7 Conclusion
In this work, we introduce ClaimDB, a fact-verification benchmark over large-scale structured data. It contains 80 unique databases, each with an average of 11 tables and 4.5M records (roughly 110M tokens when converted to text). Each claim is grounded in evidence deliberately composed of millions of records across multiple tables. ClaimDB is a clear step toward real-world claim verification—mirroring the various challenges of high-stakes fact-checking in our contemporary world.
Lastly, we conduct extensive experiments with 30 state-of-the-art open- and closed-source models, and find that more than half remain below 55% accuracy. We further uncover systematic failures around abstention—the ability to admit there is not enough evidence to decide—raising concerns about the reliability of LLMs in high-stakes verification.
@@ 8 Limitations
Reliance on BIRD. Our benchmark is derived from NL-to-SQL pairs in the BIRD benchmark (Section 3.2). As a result, any annotation error in BIRD directly propagates to ClaimDB. For example, if the NL question is “How old was the oldest mayor of Seattle when elected?” but the associated SQL query instead computes the answer for Portland, any claims generated from this pair would be incorrectly labeled. To mitigate this risk for public evaluation and model development, we construct our test splits exclusively from claims derived from the dev split of BIRD, which has undergone substantial subsequent cleanup by prior work (e.g., Wretblad et al., 2024; Liu et al., 2025). Importantly, our dependency on BIRD is not irreversible: as additional errors in BIRD are identified, the corresponding derived claims can be removed from ClaimDB on-the-fly.
Single Evidence Modality. ClaimDB deliberately focuses on structured data as its primary evidence source (which may include numerical, categorical, temporal, and short textual fields). While real-world fact verification often involves multi-modal evidence—such as free-form text, reports, news articles, charts, and other unstructured sources—structured data remains a critical yet comparatively underexplored modality in existing benchmarks. By isolating this setting, ClaimDB enables targeted evaluation of reasoning over complex structured evidence. We acknowledge that this focus alone does not capture the full breadth of real-world verification, and view extension to multi-modal evidence as a natural future direction.
Snapshot Validity and Evaluation Scope. Entailment, contradiction, and NEI are assessed with respect to the fixed database snapshot used to construct the benchmark. Because the underlying databases come from established public dataset releases—and are therefore static and potentially dated—answers may differ if a system incorporates newer external information. For example, a query about the “bottom-5 enrollment schools” could change, even if the claim was correct in the snapshot used for claim generation. As with many ML benchmarks, evaluation should be interpreted relative to the provided database state Li et al. (2023) rather than a live, continuously updated world.
@@ 9 Ethical Considerations
Data Licensing. ClaimDB is constructed from NL-to-SQL pairs derived from the BIRD benchmark, which is released under the CC BY-SA 4.0 license. Accordingly, the ClaimDB dataset is also released under the same license, in compliance with the original license terms.
@@ 10 Acknowledgments
We thank the anonymous reviewers for their constructive feedback on this paper. This work was supported by NSF III 2507117, NSF SHF 2312195, and NSF IIS 2314527. We thank Dean Light for many insightful discussions throughout the review process. We also thank the UW Database Group members for helpful feedback and David Alexander for conversations on benchmark performance.
Appendix A Tokens Per Database
In this section, we estimate the number of tokens required to represent a database in context. This analysis supports one of our central claims: in ClaimDB, the context of each claim (i.e., its underlying database) exceeds the context window of modern LLMs, which forces solutions to reason in executable programs (e.g., SQL).
To obtain this estimate, we take a straightforward approach. Each database in ClaimDB consists of multiple tables. For each database, we export all of its tables as CSV files (header row followed by all data rows), concatenate them into a single text file, and count the resulting tokens using the tokenizer of the gpt-5 series. We use this tokenizer as a representative modern LLM tokenizer.
Importantly, this estimate is conservative for two reasons. First, it does not encode the relationships and constraints present in the database (e.g., keys, foreign keys, and other schema-level structure). Second, providing structured data in context to an LLM typically requires an explicit representation—such as JSON (Singha et al. 2023)—that clearly associates each value with its column and preserves structure. This additional encoding would substantially increase the token count. The reported numbers should therefore be interpreted as a lower bound. Figure 9 shows the exact token counts for each of the 80 databases.
Figure 9: Token count per database (log scale), computed using the tokenizer of the gpt-5 series. Each bar corresponds to one of the 80 databases in ClaimDB. For more than half of the databases, naive in-context “reading” of the full evidence is practically impossible. Notably, the reported average of 110M tokens of context per claim in Table 1 reflects the fact that claims are unevenly distributed across databases. In particular, we end up with a lot more claims based on larger databases which means the per-claim token average is higher than a simple average over the 80 databases.
Appendix B Evaluation
B.1 Evaluation Setup
@@
ministral-3:3b 256K 0.15 50 90
ministral-3:8b 256K 0.15 50 90
ministral-3:14b 256K 0.15 50 90
@@
B.2 Structured Outputs
@@
For example, some leading agentic frameworks simulate structured outputs by injecting a prompt that forces the model to call a predefined tool for its final answer [^3] https://github.com/pydantic/pydantic-ai/issues/242, https://github.com/openai/openai-agents-python/issues/1778#issuecomment-3316092585 , rather than enforcing the schema at decoding time when the Grammar Constraint Decoding (GCD) option is available (e.g., on open-source models). Aside from bloating the model context, this approach is inherently brittle: the LLM can simply ignore the instruction.
@@ Appendix C Prompts
In Figures 17, 18 and 19 we showcase the prompts used in creating the claims (Section 3.5). In Figures 20 and 21 we illustrate the two prompts used for the judging process—one for contradicted and entailed claims and one for NEI since the latter has extra rubrics and extra golden context (Section 4.1). Lastly, in Figure 22, we show the single, optimized verifier prompt for all models in our evaluations.
Appendix D SQL Queries and ASTs
This section provides concrete examples of SQL queries together with their corresponding abstract syntax trees (ASTs). Figures 1316 illustrate how the filtering rules described in Section 3.3 are applied in practice.
Appendix E Similarity Distributions
SELECT business_id
FROM Business
WHERE state LIKE AZ AND stars = 5
Figure 13: ✗ Not Compositional. The SQL query (top) and its AST (bottom). The query is excluded since it only applies local filters (state, stars) and does not involve aggregations, joins, or other operations that combine information across many records.
SELECT COUNT(T1.inspection_id)
FROM inspection AS T1 INNER JOIN employee AS T2
ON T1.employee_id = T2.employee_id
WHERE T2.first_name = Lisa AND
T2.last_name = Tillman AND
T1.results = Out of Business
Figure 14: ✓ Compositional. The SQL query (top) and its AST (bottom). The query is retained since its AST contains an aggregate function (COUNT), which counts a large number of records and collapses them into a single value (the count). Verifying a claim derived from this query therefore requires compositional reasoning over many records, which cannot be performed by naively “reading” the records in-context and counting by an LLM alone.
SELECT T2.movie_title, T1.user_id,
T1.rating_score, T1.critic
FROM ratings AS T1 INNER JOIN movies AS T2
ON T1.movie_id = T2.movie_id
WHERE T1.critic IS NOT NULL
Figure 15: ✗ Not Compositional. The SQL query (top) and its AST (bottom). The query is excluded since it does not contain aggregations, orderings, window functions, or joins involving three or more tables (see Section 3.3).
SELECT T1.title
FROM movie AS T1
INNER JOIN movie_keywords AS T2
ON T1.movie_id = T2.movie_id
INNER JOIN keyword AS T3
ON T2.keyword_id = T3.keyword_id
WHERE T3.keyword_name = extremis
Figure 16: ✓ Compositional. The SQL query (top) and its AST (bottom). The query is retained since it joins three tables together—the tables movie aliased as T1, movie_keywords aliased as T2, and keyword aliased as T3—which is a compositional operation as defined in Section $3.$ A verifier must explicitly combine these three tables in order to verify a claim derived from this SQL query.
@@ Appendix E Similarity Distributions
Figure 22: Prompt for all verifier models in the experiments (Section 6.1). It has been carefully constructed with many iterations and refinements across the different model families. We have tried our best to make the models perform as good as possible. For example, after noticing some models deciding not to use tools and hallucinate an answer, we including instructions like “Use the available tools to query the database and gather evidence before making a decision” and “You should always start by querying the database for the schema”. Furthermore, for a discussion on output format see Appendix B.2. Notably, for the observant readers, there is no explicit explanation of the tool the agent has in its disposal because these descriptions are provided by the frameworks internally (most are built this way) by injecting the docstrings and input-output schemas of the tools in the prompt using templates.